# backgroundcolor=gray!3, roundcorner=3pt

### v1

@@ 5.2.2 Performance of guidance with human advisor
Fig. 8 shows the cumulative rewards of the agent with a human advisor under two different advice quotas.
Key performance improvement trends are retained. We see similar trends to those in Sec. 5.2.1. That is, advised agents outperform the unadvised agent even with a severely reduced advice quota. As a reminder, the “all” strategy provided advice about every cell; the “holes&goal” strategy provided advice about the holes and the goal, i.e., about 20% of the cells—here, the human operating under 10% and 5% quotas is still able to improve the agent’s performance. With 10% quota, even at $u=0.4$, the performance reaches that of the synthetic advisor.
The role of certainty increases as the advice quota is reduced. This trend is indicated by the distances between the lines both in the linear and log scales between 10% and 5%. At 10% advice quota, the performance differences between different levels of certainty are similar; at 5% advice quota, performance drops more prominently as certainty decreases.
Advice quotas: differences more pronounced in low-certainty cases. The difference between the two quotas is more obvious as certainty decreases. In the 5% case, the fully certain advice $u=0.0$ only reaches the level of somewhat certain advice $u=0.2$ of the 10% case. Above uncertainty level $u=0.2$, the 5% quota deteriorates substantially, marking the likely limit at which opinion-based guidance can be efficient.
@@
Key improvements are retained. We see similar trends to those in Sec. 5.2.1. The policy in advised agents is better reinforced, as indicated by the darker shades of blue, especially around holes and the goal. We see more explored states by advised agents, though the difference between moderately certain advice with low advice quota, and no advice is less substantial.
Certainty becomes more of a factor. In line with the observations about the cumulative reward in the human-advised mode, certainty becomes an important factor in advice. The policy under 5% advice quota and at $u=0.4$ shows minor improvements over the unadvised mode, while fully certain advice at $u=1.0$ still shows substantial differences.
@@
[…]
@@ 6.1 The impact of advice – Discussion of the results
Even uncertain advice is useful
The key takeaway of this work is that even uncertain opinions improve the performance of the learning agent. Even with a human advisor, opinions as uncertain as $u=0.6$ guided agents outperform vanilla agents. Considering that opinions are formed much faster than hard evidence can be produced, opinion-based guidance can be an important improvement over traditional guidance mechanisms if opinions can be efficiently elicited from advisors.
@@
To provide an intuition of what uncertainty at the level of $u=0.6$ means, e.g., in terms of the Frozen Lake problem, consider the two edge cases when an advisor is very sure about a state being (1) disadvantageous and (2) advantageous. Given that $u=0.6$, the remaining opinion weight is $0.4$, which has to be distributed over disbelief d and belief b. The first case would translate to an opinion $\omega_{1}=(0.4,0.0,0.6,0.25)$, and the second to $\omega_{2}=(0.0,0.4,0.6,0.25)$. Using the formulas from Sec. 3.2, the two opinions translate to $P_{1}=0.0+0.15=0.15$ and $P_{2}=0.4+0.15=0.55$, respectively. That is, the advisor has between 0.15–0.55 probability—good for a meager 40% wiggle room—to express the advantage of choosing a particular state. The advisor cannot express anything below 15% and above 55%, and yet, these values are still vastly superior to not giving advice without misleading the agent.
Finding trade-offs between certainty, exhaustiveness, and cost of advice
A key trend we observe is that performance deteriorates as uncertainty increases. In some cases, low-certainty advice might have a negative impact on the performance. Shown in Tab. 10, agents advised at $u=1.0$ perform below the vanilla agent, although Figures 5(c)-5(d) and 7(h)-7(i) show a narrowing gap between low-certainty advice and no-advice agents, indicating that low-certainty advice would likely catch up with no-advice agents. Thus, the safe level of uncertainty in our experiments is around $u=0.6$, and any advice beyond that level of uncertainty is not much of a help. Adopters of the approach should assess the limits of their own settings and decide about the accepted levels of uncertainty accordingly.
An interesting trend to observe with the synthetic advisor is that the “holes&goal” strategy reinforces some key cells better than the “all” strategy, e.g., around the goal and around the starting point. At the same time, the “holes&goal” strategy allows to explore more states at higher levels of certainty than the “all” strategy (cf. Fig. 7(e) and Fig. 7(c)); however, this relation turns around in the more uncertain cases (cf. Fig. 7(b) and Fig. 7(d)). It is plausible to conclude that selective and targeted advice strategies might lead to more pronounced differences in the shaped policy, and by allowing for an improved exploration rate, they transitively allow for a better-reinforced policy. Thus, adopters of our approach should look for trade-offs between exhaustiveness and cost of advice as the returns of facilitating advising the agents might diminish.
Ultimately, adopters should seek trade-offs between each of these quality properties: certainty, exhaustiveness of advice, and cost of advice. We foresee the identification and management of these trade-offs to become the key to the successful and practical adoption of our approach.
Synthetic vs. Human advice
Comparing synthetic and human advisors’ performance does not show substantial differences (Tab. 10). In particular, the human advisor with sufficient quota achieves performance similar to that of the synthetic advisor. In our experiments, the human with 10% advice quota (i.e., the number of pieces of advice is 10% of the number of all cells) achieved similar scores at $u=0.0$$u=0.6$ as the synthetic advisors that had 100% (“all”) and 20% (“holes&goal”) quota. This suggests that human creativity is superior to synthetic strategies, both to exhaustive style and strategically targeted ones. Thus, ours is another one in the long line of works that highlight the role of human cognition in computer-automated methods, such as machine learning. Since the human advisor can quickly become the bottleneck in reinforcement learning, adopters of our approach should consider implementing efficient advice strategies, particularly asynchronous ones in which human and machine agents can work independently.
@@ 6.2 Open challenges and research opportunities
In our approach, we opted for a policy-based RL algorithm because it allows for better investigation of the impact of opinion on the explicitly represented policy. We are fairly confident that our method translates to value-based techniques as well, but validation is left for future work. A fairly more complicated challenge is introducing guidance into deep reinforcement learning, where the policy is encoded in a neural network, and the lack of explainability might hinder the development of a sound guiding framework. With the demonstrated utility of deep reinforcement learning in numerous problems, such as the management of engineering systems [3] and simulator inference [19], its support by opinion-based guidance seems like a natural next step.
Interaction patterns for optimal guidance
We relied on one-time advice provided before exploration. However, more sophisticated advice dynamics have been used in guided reinforcement learning [41, 45]. We recommend researching more interactive patterns of advising. A known drawback of placing humans into the learning loop is the high cost of human reward signals, which are caused by the human stakeholders quickly becoming performance bottlenecks [33]. We see opportunities in opinion-based guidance in this aspect as opinions might be easier to formulate than hard evidence, alleviating the load on the human. Another interesting research opportunity our approach motivates is the investigation of push and pull mechanisms in advising reinforcement learning agents, in which the human can push advice toward the agent, and the agent can actively query the human for advice at any time. Such mechanisms would allow for the human to provide advice under limited quotas when advice matters the most.
Languages for expressing advice
A key advantage of opinion-based guidance is the rapid pace at which opinions emerge. To leverage this advantage to its fullest, intuitive languages are needed to express opinions. Domain-specific languages (DLSs) [46] have been widely used to capture expert opinions at high levels of abstraction, narrowing the gap between the modeling language and human cognition. However, DSLs typically operate with reduced syntax to remain truly specific to the problem domain, and therefore, automated methods for DSL engineering are in high demand [9]. This is true for guided reinforcement learning as well. We envision language workbenches that allow for the automated construction of DSLs through which advice can be provided by domain concepts. For example, in the Frozen Lake example, one might want to use descriptive quantifiers, such as “hazardous” or “beneficial” instead of relying on an integer scale. Unfortunately, there is little research on domain-specific languages for reinforcement learning, and existing languages mostly focus on programming reinforcement learning procedures [40] rather than expressing opinions at high cognitive levels. We see plenty of room for research targeting domain-specific modeling languages for reinforcement learning, inspired by the advancements in the model-driven engineering community [18, 35, 54].
@@ Appendix A Study setup
1[11,11], +2
2[0,4], -2
3[0,5], -2
4[1,4], -2
5[1,5], -2
6[1,7], -2
7[1,8], -2
8[1,9], -2
9[1,11], -2
10[2,7], -2
11[3,0], -2
12[3,7], -2
13[4,4], -2
14[4,9], -2
15[5,5], -2
16[6,3], -2
17[6,4], -2
18[6,5], -2
19[6,7], -2
20[7,1], -2
21[7,3], -2
22[8,3], -2
23[9,1], -2
24[9,4], -2
25[9,6], -2
26[10,1], -2
27[10,4], -2
28[11,10], -2
Listing 3: Input advice with an opinion about every cell of Fig. 10. Corresponds to the HOLES&GOAL strategy in the experiments.
@@
Listing 4: Input advice with an opinion about every cell of Fig. 10. Corresponds to the ALL strategy in the experiments.

### v2

@@ 5.2.2 Performance of guidance by a single human advisor
Fig. 9 shows the cumulative rewards of the agent with a human advisor under the two advice quotas.
@@
Key performance improvement trends are retained
We see similar trends to those in Sec. 5.2.1. That is, advised agents outperform the unadvised agent even with a severely reduced advice quota. The human operating with 10% and 5% quotas is still able to improve the agent’s performance. With 10% quota, the performance typically reaches that of the oracle. Moreover, at $u=0.2$ and $u=0.8$, the human with 10% quota performs better than any oracle. At $u=0.4$ and $u=0.6$, the human advisor with 10% advice quota performs better than the oracle with 20% advice quota. The only case in which the human does not outperform at least one of the oracles is $u=0.0$; but even in this case, the human’s performance is only about 5% below of that of the oracles.
The role of certainty increases as the advice quota is reduced
This trend is indicated by the distances between the lines both in the linear and log scales between 10% and 5%. At 10% advice quota, the performance differences between different levels of certainty are similar; at 5% advice quota, performance drops more prominently as certainty decreases.
Advice quotas: differences more pronounced in low-certainty cases
The difference between the two quotas is more obvious as certainty decreases. In the 5% case, even the fully certain advice $u=0.0$ performs worse than the somewhat certain advice of $u=0.2$ in the 10% case. Starting at $u=0.2$, the 5% quota deteriorates substantially; but still outperforming the unadvised agent.
@@
Key improvements are retained
We see similar trends to those in Sec. 5.2.1. The policy in advised agents is better reinforced, as indicated by the darker shades of blue, especially around holes and the goal. We see more explored states by advised agents, though the difference between moderately certain advice with low advice quota, and no advice is less substantial.
Certainty becomes more of a factor
In line with the observations about the cumulative reward in the human-advised mode, certainty becomes an important factor in advice. The policy under 5% advice quota and at $u=0.4$ shows minor improvements over the unadvised mode, while fully certain advice at $u=0.0$ still shows substantial differences.
@@
5.2.3 Performance of guidance by cooperating human advisors with partial information
Advisor type
In this experiment, we use the advice of two human advisors who have partial knowledge about the problem. Partial information is modeled by the advisors not giving advice about cells that are beyond a given distance, as the increasing uncertainty would discount these pieces of advice, rendering their effect on the policy virtually nil.
Advice quota
We evaluate two advice quotas. In the first case, both human advisors provide advice about 10% of the cells at their own discretion. In the second case, both human advisors provide advice about 5% of the cells at their own discretion.
Uncertainty calibration
As opposed to the previous experiments, where we evaluated guidance at multiple levels of uncertainty, here, we calibrate uncertainty for each advice, based on the distance between the advisor and the location of the advice.
Cooperative advice type
We evaluate two types of cooperation: sequential cooperation (one human in the top-left corner and on in the bottom-right corner), and parallel cooperation (one human in the top-right corner and one in the bottom-left corner).
Cumulative reward
Fig. 11 shows the cumulative reward of the agent with two cooperating advisors under two different advice quotas.
Linear scale
(f) Advice quota: $10\%$
(g) Advice quota: $5\%$
Log scale
(h) Advice quota: $10\%$
(i) Advice quota: $5\%$
Figure 11: Cumulative reward with advice from two cooperating human advisors
Performance improvement over the unadvised agent is retained
Similar to the previous experiments, advised agents outperform the unadvised agent, despite the severely reduced advice quota, and the advisor’s partial information about the problem space.
Sequential cooperation outperforms parallel cooperation
This performance difference is consistent both with 10% and 5% advice quota. Parallel cooperation at 10% shows similar performance to sequential cooperation at 5%.
Cooperation with partial information can perform comparably to fully informed humans
[…]
Learning dynamics
Learning dynamics are affected in both cooperative cases, but the performance deterioration is more pronounced in the 5% case. The learning curve is both delayed and less steep as compared to the previous experiments. Still, the pace of learning is significantly more rapid than that of the unadvised agent.
Effects on the policy
Cooperative advice outperforms single human advice
Specifically, sequential cooperation results in a better-reinforced policy and allows for more explored states in more uncertain situations. This is can be inferred from the difference between Fig. 12 and Fig. 10: we encounter more darker cells (better reinforced local decisions) and less empty cells (less unexplored states). The effect on the policy is weaker to that of the oracle.
Sequential cooperation is better in reinforcement and parallel cooperation is better in exploration performance
In pattern unique to cooperative guidance, we observe that heatmaps visualizing sequential cooperation show stronger reinforced local decisions (Fig. 12(b) over Fig. 12(c), and Fig. 12(d) over Fig. 12(e)); while heatmaps visualizing parallel cooperation show better coverage of the problem (Fig. 12(c) over Fig. 12(b), and Fig. 12(e) over Fig. 12(d)). In fact, parallel cooperative guidance at 10% (Fig. 12(c)) explores as many states as the oracle at 100% quota and $u=0.0$ uncertainty (Fig. 8(c)), i.e., the fully informed and fully certain oracle.
(a) No advice
Advice quota: 10%
(b) Sequential
(c) Parallel
Advice quota: 5%
(d) Sequential
(e) Parallel
Figure 12: Heatmap visualization of the final learned policy with two human advisors
@@ 5.2.3 Performance of guidance by cooperating human advisors with partial information
frametitle= RQ2 , innertopmargin=2pt, frametitleaboveskip=-frametitlealignment= {mdframed} Our observations indicate that opinion-based guidance contributes to significant performance improvement in reinforcement learning agents, even at moderate-to-high uncertainty. Human advisors, even with substantially lower advice quotas, can perform at the level of an oracle, and might even outperform them. Uncertainty becomes more of a factor as advice quota decreases. In realistic settings of cooperating humans with partial and complementary information about the problem, we find that guidance, even at low advice quotas, still results in significant improvements over unadvised agents.
@@ 6.1 The impact of advice – Discussion of the results
Even uncertain advice is useful for guidance
The key takeaway of this work is that even uncertain opinions improve the performance of the RL agent. The only case in which an opinion-guided RL agent did not outperform the unadvised agent is in the case with a single human advisor with 5% quota at uncertainty level $u=0.8$ (Fig. 8(g)).
@@
To provide an intuition of what uncertainty at the level of $u=0.8$ means, e.g., in terms of the Frozen Lake problem, consider the two edge cases when an advisor is sure about a state being (i) disadvantageous and (ii) advantageous. Given that $u=0.8$, the remaining opinion weight is $0.2$, which has to be distributed over disbelief d and belief b. Case (i) translates to an opinion $\omega_{i}=(0.2,0.0,0.8,0.25)$, and the second to $\omega_{ii}=(0.0,0.2,0.8,0.25)$. Using the formulas from Sec. 3.2, the two opinions translate to $P_{i}=0.2+0.25\times 0.8=0.6$ and $P_{ii}=0.0+0.25\times 0.8=0.4$, respectively. That is, the advisor can express they believe the probability of a particular state being advantageous should be between 0.4–0.6 probability—good for a meager 20% wiggle room. In other words, the advisor cannot express anything below 40% and above 60% which, evidently, is the descriptive power an unadvised agent has with 5% advice quota (Fig. 8(g)). (At higher advice quotas, the same 0.4–0.6 advice range, i.e., the same uncertainty of $u=0.8$ is used more efficiently by the advisors, resulting in performance improvements over the unadvised agent.)
Considering that opinions emerge earlier than hard evidence can be produced [18], opinion-based guidance positions to be an important improvement over traditional guidance mechanisms. To leverage the benefits of opinion-based guidance, opinions should be elicited efficiently (Sec. 6.2).
Effectiveness of guidance is impacted by certainty and exhaustiveness
Two key trends we observe are that performance deteriorates as uncertainty increases and that performance deteriorates as advice quota decreases. We observe these trends in each of our experiments (see Tab. 10 and Fig. 6). The performance drop between advice quotas is more pronounced at higher levels of uncertainty. Specifically, at $u=0.8$, advice loses its guidance value at 5% quota as the advised agent does not perform better than the unadvised one (single human case, Fig. 8(g)). This combination of parameters seems to be the lower bound of utility of our approach.
Another interesting trend to observe is that at high levels of certainty ($u=0.0$) the human advisor with 5% advice quota reinforces as many key decisions as with 10% quota (cf. Fig. 10(e) and Fig. 10(c)). Furthermore, two cooperating human advisors, each with 5% advice quota, reinforce the agent’s policy better than a single advisor with 10% quota (cf., e.g., Fig. 12(d) and Fig. 10(c)).
Thus, while higher certainty and higher advice quota (exhaustiveness) improve the effectiveness of guidance, the cost of a thorough advice strategy might not be always justified. We recommend adopters of our approach to seek trade-offs between certainty, exhaustiveness, and the costs of providing advice.
Human advisors are as effective as idealized oracle advisors
Comparing the performance of the oracles and human advisors (Tab. 10), we do not observe substantial differences. The single human advisor with sufficient quota achieves a performance similar to that of the oracle. The performance is within 5% in three of five evaluated cases between oracles and human advisors ($u=0.0,0.2,0.6$), with one of these cases showing an improvement of the human over the oracle ($u=0.2$). In two of five evaluated cases, the difference is more pronounced, once in favor of the oracle ($u=0.4$) and once in favor of the human advisor ($u=0.8$). Improvements due to human advice are particularly apparent in highly uncertain situations.
We attribute these observations to human creativity in providing advice. Thus, ours is another one in the long line of works that highlight the role of human cognition in computer-automated methods, such as machine learning [31, 5].
Since the human advisor can quickly become the bottleneck in RL, adopters of our approach should consider implementing efficient advice strategies, particularly asynchronous ones in which human and computer-automated agents can work independently.
Cooperative guidance with partial information performs well compared to advisors with complete information
Our experiments show that the performance of cooperative guidance with partial information about the problem is only slightly below the performance of oracles and single human advisors with complete information about the problem. At $u=0.4$, sequential cooperation with $2\times 5$% quota outperforms the comparable single human advisor with 10% quota (5 544.100 vs 4 727.433; +17.28%) and the oracle with twice as much quota of 20% (5 544.100 vs 4 218.000; +31.44%). Related to this latter comparison, at $u=0.2$ sequential cooperation with $2\times 10$% quota performs close to the oracle with comparable quota (20%) – 7 551.567 vs 7 835.267 (-3.62%). These results demonstrate that partial information equates to moderate levels of uncertainty in fully-informed situations. The cooperative experiments we conducted emulate a highly realistic setting. Thus, it is plausible to expect that the benefits of opinion-based guidance can be retained in real applications.
Comparing cooperation modes, we observe that sequential cooperation gains more (Fig. 10(f) and Fig. 10(g)) cumulative reward and earlier (Fig. 10(h) and Fig. 10(i)) than parallel cooperation. In sequential cooperative guidance, uncertainty is lower in critical areas of map, i.e., around start, goal, and along path (Fig. 12(b) vs Fig. 12(c); and Fig. 12(d) vs Fig. 12(e)). This means advice about critical areas has more of an impact on the agent’s policy, which leads to higher cumulative reward. In particular, having an advisor located at the agent’s start state where uncertainty is low helps the agent navigate the early parts of the map better. Thus, the agent is able to avoid holes early and take more steps in early episodes (more exploration), leading to learning the optimal policy earlier (as demonstrated in the log charts Fig. 10(h) and Fig. 10(i)).
@@ 6.2 Open challenges and research opportunities
Interaction patterns for optimal guidance
We relied on one-time advice provided before exploration. However, more sophisticated advice dynamics have been used in guided RL [46, 51]. We recommend researching more interactive patterns of advising. A known drawback of placing humans into the learning loop is the high cost of human reward signals, which are caused by the human stakeholders quickly becoming performance bottlenecks [38]. We see opportunities in opinion-based guidance in this aspect as opinions might be easier to formulate than hard evidence, alleviating the load on the human. Another interesting research opportunity our approach motivates is the investigation of push and pull mechanisms in advising RL agents, in which the human can push advice toward the agent, and the agent can actively query the human for advice at any time. Such mechanisms would allow for human advisors to provide advice under limited quotas when advice matters the most.
Languages for expressing advice
A key advantage of opinion-based guidance is the rapid pace at which opinions emerge. To leverage this advantage to its fullest, intuitive languages are needed to express opinions. Domain-specific languages (DLSs) [52] have been widely used to capture expert opinions at high levels of abstraction, narrowing the gap between the modeling language and human cognition. However, DSLs typically operate with reduced syntax to remain truly specific to the problem domain, and therefore, automated methods for DSL engineering are in high demand [10]. This is true for guided RL as well. We envision language workbenches that allow for the automated construction of DSLs through which advice can be provided by domain concepts. For example, in the Frozen Lake example, one might want to use descriptive quantifiers, such as “hazardous” or “beneficial” instead of relying on an integer scale. Unfortunately, there is little research on domain-specific languages for RL, and existing languages mostly focus on programming RL procedures [45] rather than expressing opinions at high cognitive levels. We see plenty of room for research targeting domain-specific modeling languages for RL, inspired by the advancements in the model-driven engineering community [20, 40, 60].
@@
In our approach, we opted for a policy-based RL algorithm because it allows for a better investigation of the impact of opinion on the explicitly represented policy. We are fairly confident that our method translates to value-based techniques as well, but validation is left for future work. A fairly more complicated challenge is introducing guidance into deep reinforcement learning, where the policy is encoded in a neural network, and the lack of explainability might hinder the development of a sound guiding framework. With the demonstrated utility of deep reinforcement learning in numerous problems, such as the management of engineering systems [3] and simulator inference [21], its support by opinion-based guidance seems like a natural next step.
7 Threats to validity and limitations
Construct validity
Observations of improved exploration performance may be artifacts of fortunately aligned problems (here: maps) or a few fortunate pieces of advice rather than artifacts of systematic effects. To mitigate this threat to validity, first, we experimented with different map layouts and observed the same performance trends. To facilitate consistent manual experimentation, we developed a map generator that generates map layouts in a reproducible way. Second, we ran our scaled-up experiments with a variety of settings, including both human and oracle advisors. A negligible threat to validity in cooperative experiments stems from the human advisors’ strategy of considering their own level of certainty when giving advice. Human advisors prefer not to waste their limited advice quota and thus, will be likely biased towards giving advice with higher certainty. Thus, the positive results of cooperative guidance are artifacts of opinion-based guidance and human behavior. For the purposes of our research questions, human behavior does not threaten validity.
Internal validity
Evaluating the approach with human advisors inevitably poses threats to internal validity. Specifically, maturation and testing might have influenced our results as the authors experimented with advising strategies and advice values. For example, cells that are not holes or goals are labeled freely with -1, 0, or 1, and a human might learn how to misuse these values to achieve a specific behavior by the agent. We mitigated these threats by defining clear advise score rules, which are also used by the oracle advisor.
External validity
We chose only one flavor of RL—a policy-based method—to demonstrate our approach, and we did not provide evidence that our approach can be applied to other RL methods. Although we are reasonably confident that our methods translate to RL methods where the policy is not explicitly represented (e.g., value-based methods), future work will investigate this question in more detail.
Limitations
We used a relatively simple guidance strategy, in which advice is provided once before the exploration. It is reasonable to assume that more interactive guidance modes might shed light on additional benefits and challenges. However, we are reasonably confident that the scope of this article was sufficient to draw key takeaways that hold in more interactive reinforcement learning setups as well. Nonetheless, we will experiment with interactive guidance modes in future work.
@@ Appendix A Study setup
Appendix B Input artifacts
B.1 Oracle
B.1.1 Advice – 100% quota, oracle
@@
Listing 3: Advice of the oracle with 100% quota, with an opinion about every cell of Fig. 13.
B.1.2 Advice – 20% quota, oracle
1[11,11], +2
2[0,4], -2
3[0,5], -2
4[1,4], -2
5[1,5], -2
6[1,7], -2
7[1,8], -2
8[1,9], -2
9[1,11], -2
10[2,7], -2
11[3,0], -2
12[3,7], -2
13[4,4], -2
14[4,9], -2
15[5,5], -2
16[6,3], -2
17[6,4], -2
18[6,5], -2
19[6,7], -2
20[7,1], -2
21[7,3], -2
22[8,3], -2
23[9,1], -2
24[9,4], -2
25[9,6], -2
26[10,1], -2
27[10,4], -2
28[11,10], -2
Listing 4: Advice of the oracle with 20% quota, with an opinion about the holes and the goal cell of Fig. 13.
B.2 Single human experiments
B.2.1 Advice – 10% quota, single human advisor
1[1,1], +1
2[1,4], -2
3[2,2], +1
4[3,4], +1
5[3,5], +2
6[3,7], -2
7[4,4], -2
8[4,6], +2
9[4,9], -2
10[5,5], -2
11[5,7], +2
12[6,7], -2
13[11,10], -2
14[11,11], +2
Figure 14: Advice of the single human advisor with $10\%$ quota, and its visualization.
B.2.2 Advice – 5% quota, single human advisor
1[1,4], -2
2[3,7], -2
3[4,4], -2
4[5,5], -2
5[6,7], -2
6[11,10], -2
7[11,11], +2
Figure 15: Advice of the single human advisor with $5\%$ quota, and its visualization.
B.3 Cooperative human experiments
B.3.1 Sequential cooperation, 10% quota each
1[0,4], -2
2[0,5], -2
3[1,4], -2
4[1,5], -2
5[1,6], -1
6[2,4], +1
7[2,7], -2
8[3,0], -2
9[3,4], +1
10[3,5], +1
11[3,7], -2
12[4,4], -2
13[4,6], +1
14[5,5], -2
Figure 16: Advice from the top-left corner in the sequential cooperative human experiments, with $10\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 17.
1[1,10], -1
2[1,11], -2
3[4,9], -2
4[5,7], +1
5[5,8], +1
6[5,9], +1
7[6,5], -2
8[6,6], -1
9[6,7], -2
10[9,4], -2
11[9,6], -2
12[10,4], -2
13[11,10], -2
14[11,11], +2
Figure 17: Advice from the bottom-right corner in the sequential cooperative human experiments, with $10\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 16.
B.3.2 Sequential cooperation, 5% quota each
1[1,4], -2
2[1,5], -2
3[1,6], -1
4[2,7], -2
5[3,7], -2
6[4,4], -2
7[5,5], -2
Figure 18: Advice from the top-left corner in the sequential cooperative human experiments, with $5\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 19.
1[4,9], -2
2[5,7], +1
3[6,5], -2
4[6,7], -2
5[9,4], -2
6[11,10], -2
7[11,11], +2
Figure 19: Advice from the bottom-right corner in the sequential cooperative human experiments, with $5\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 18.
B.3.3 Parallel cooperation, 10% quota each
1[1,4], -2
2[1,5], -2
3[1,6], -1
4[1,7], -2
5[1,8], -2
6[1,9], -2
7[1,10], -1
8[1,11], -2
9[2,7], -2
10[3,7], -2
11[4,9], -2
12[6,7], -2
13[10,11], +1
14[11,11], +2
Figure 20: Advice from the top-right corner in the parallel cooperative human experiments, with $10\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 21.
1[4,4], -2
2[5,5], -2
3[6,3], -2
4[6,5], -2
5[7,1], -2
6[7,3], -2
7[8,3], -2
8[9,1], -2
9[9,4], -2
10[9,6], -2
11[10,1], -2
12[10,4], -2
13[11,10], -2
14[11,11], +2
Figure 21: Advice from the bottom-left corner in the parallel cooperative human experiments, with $10\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 20.
B.3.4 Parallel cooperation, 5% quota each
1[1,4], -2
2[2,7], -2
3[3,7], -2
4[4,9], -2
5[6,7], -2
6[10,11], +1
7[11,11], +2
Figure 22: Advice from the top-right corner in the parallel cooperative human experiments, with $5\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 23.
1[4,4], -2
2[5,5], -2
3[6,3], -2
4[6,5], -2
5[7,3], -2
6[9,4], -2
7[11,10], -2
Figure 23: Advice from the bottom-left corner in the parallel cooperative human experiments, with $5\%$ quota, and its visualization. The counterpart of the cooperative advice is shown in Fig. 22.